Shift Summary for owl shift, 25/sep/99

AG (ag@fnal.gov)
Sat, 25 Sep 1999 07:58:52 -0400 (EDT)

Shift Summary for owl shift, 25/sep/99

On Shift: SG, AG

Hours of beam available at all intensities: 2h 25min
Hours of beam available at nominal intensity: 2h 25min
Hours of E799 data (special runs not included): 2h 25min
Hours of special runs (laser scans, muon runs...): 0h 0min

Accounting of lost time:

TeV abort + NM2EU power supply failure.

========================================================================

Runs ended this shift:

Run 14714 - 6.3M - BEAM_799

Tapes brought to FCC this shift:
(including tapes for online split)

None.

========================================================================

Accesses during shift:

None.

========================================================================

Overview of shift:

Stable beam and smooth e799 data taking at low (~7.6e12) intensity till
beam dies at 02:36. At 05:00 MCR calls to tell that we have serious
problem with NM2EU power supply. They call Rick C.
Looks like the power supply has to be replaced.
There will be no beam in KTeV for ~8-10 hours.

========================================================================

Description of Shift:

run____spill______time_______comment

---------------------------------------------------------

14714____________0010____Beam is back.

_________________0135______Spills 152-156 are bad -> laser crate
____________________________DYC full, after kfrend got ktevana errors and
____________________________pipe-line errors -> reset pipeline+kfrend

_________________0236______TeV abort, NM2EU tripped.

_________________0500______MCR calls to tell that they have serious
____________________________problem with NM2EU power supply, and that they
____________________________will call Rick C. about this. They say there
____________________________will be no estimate at this time, but it's
____________________________going to be quite a while till we get beam...

14714____________0504______Stop the run and

14715____1_______0512___Start COSMICS run.

_________________0600______We notice that NM4AN tripped. We call MCR to
____________________________find out if that's connected to them working
____________________________on the power supply problem. They think it's
____________________________probably due to the experts working
____________________________and promise to call us later if
____________________________they know more... Yes, they turned down all
____________________________the magnets to replace NM2EU power supply.

_________________0623______Beam to users, but ... they do not mean KTeV
____________________________this time. Our problem is going to take ~8-10
____________________________hours to fix :-(

========================================================================

Follow up Notes on Previous Shift

========================================================================

Items to Follow up on Next Shift